The length of a rectangular garden is 9 feet longer than its width. The garden's perimeter is 190 feet. Find the length of the garden.
Answer by ewatrrr(24785)

 
Hi,
.
2w + 2L = P
2w + 2(w+9) = 190ft
4w = 172ft
w = 43ft and L = 52ft
Checking..
2w + 2L = P
86ft + 104ft = 190ft CHECKS!